# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v2

@@ B.3 Incorporating the influence of baryons
where $\rho_{s}$, and $r_{s}$ are the NFW scale parameters and $\rho_{H}$, and $r_{H}$ are the Hernquist scale parameters. The dimensionless quantities are defined as $\hat{\rho}_{H}=\rho_{H}/\rho_{s}$, $\hat{r}_{H}=r_{H}/r_{s}$, and $\hat{r}_{\rm eff}=r_{\rm eff}/r_{s}=(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{3}/2)/(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{2}/2)$. To obtain predictions that incorporate baryonic effects, we simply multiply this factor by the core size $r_{c}(\tau)$ in the parametric model for DM-only halos. While schematic, this method accounts for the baryon-induced contraction effect, which scales with the stellar-to-halo mass ratio, and smoothly reduces to the no-baryon limit as $\tau$ approaches zero.
In the main text, we have illustrate the simulation results for a median concentration halo in Fig. 3 middle panel, where baryons have growing cores along with the dark matter. The parametric model predictions are tested against the simulation results at different times. Here we provide another example for a more compact halo. In the right panel of Fig. 8, we compare the parametric model predictions with the simulation results for a $M^{11}~\rm M_{\odot}$ halo ($1\sigma$ concentration). The profiles, shown in progressively darker colors, correspond to $t=0,3,6,$ and $9$ Gyr, respectively. While small deviations in shape are present, the model effectively captures the growth of inner densities in the baryonic case (solid lines) compared to the DM-only cases (dashed lines). The baryon profiles themselves (dotted lines) remain largely unchanged in this benchmark.
Appendix C B. Calculation of strong lensing effects
Figure 9: Density profiles of cluster subhalos with $M>10^{11}\,M_{\odot}/h$ in CDM, SIDM2c, and SIDM2c+baryons from parametric model predictions (dashed). Solid curves show corresponding fits using the PSIDM model of Refs. Hou et al. (2025). Subpanels display the relative deviation between simulation and fit, defined as $2\,(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The parametric model accurately captures the density structure of simulated subhalos across a broad mass range.
@@
Specifically, the parametric SIDM (PSIDM) density profile introduced in Ref. Hou et al. (2025) has the same parametric form as Eq. 3. The evolution of its halo parameters is governed by the functions given in that reference. For CDM halos, a single PSIDM profile is sufficient for an accurate fit, as it reduces to the NFW profile in the limit $\tau\to 0$. For SIDM2c halos, we use a superposition of two PSIDM profiles to achieve an accurate representation.
Figure 9 shows the fitting performance for cluster subhalos in three models: from left to right, CDM (blue), SIDM2c (red), and SIDM2c+baryons (magenta). Only halos with masses above $10^{11}~\rm M_{\odot}$ are shown. The lower panels illustrate the relative difference between the simulated (Sim) and fitted (Fit) profiles, computed as $2(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. Clearly, the two PSIDM model reproduces the simulated halo profiles with high accuracy, even in the presence of baryonic effects.
@@
Appendix D C. Cluster density profiles and constraint
Figure 10: Left: Density profiles of cluster host halos in the CDM, SIDM2v, and SIDMx simulations (dashed). The profiles are fitted by two PSIDM density profiles (solid) Hou et al. (2025). Right: Comparison of cluster density profiles at $t=5~\rm Gyr$ for the two-component SIDM2v model (solid) and the one-component SIDM01 model (dashed). The initial CDM halo (black) and baryons (orange) are also shown. The inner densities for SIDM01 are consistently shallower than those for SIDM2v, both with and without baryons, indicating that SIDM2v remains consistent with cluster-scale constraints.
The reconstruction of cluster mass profiles has presented one of the most robust constraints on SIDM in clusters, which varies but is generically around $0.1~\rm cm^{2}/g$ Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). Core formation in clusters is efficient and only moderately influenced by baryon-induced contraction. For a constant, small self-interaction cross section per unit mass, the rate of gravothermal evolution, quantified by the inverse of the core-collapse time, is proportional to $\rho_{s}^{3/2}r_{s}$, implying that more massive halos evolve more rapidly. Moreover, the stellar mass in clusters generally constitutes only about $0.05\%$ of the total halo mass, and thus contributes negligibly to contraction.
@@
Figure 10 (left) presents the density profiles of host halos from cosmological simulations. As a result of the merging environment and its interplay with inter- and intra-species scattering, we observe intriguing twists in the density profiles of both SIDM2v and SIDMx, which complicates the assessment of cluster-scale constraints.
To obtain a robust assessment, we perform isolated simulations for the cluster halo under SIDM2v, using the scale parameters of the CDM cluster halo as a reference. We consider two contrasting scenarios with and without baryons. The baryonic component, with a mass of $7.47\times 10^{11}~\rm M_{\odot}$, is modeled as a Hernquist profile with $r_{H}=34.1~\rm kpc$. These parameters align with the median stellar-to-halo mass relation and median size-mass relation Moster et al. (2013); Carleton et al. (2019). We use $5\times 10^{6}$ dark matter simulation particles in both cases to better resolve the inner halo regions. Baryons are treated as collisionless stars with particle masses equal to those of the dark matter.
Figure 10 (right) shows the simulation results (solid) at $t=5~\rm Gyr$, compared with the one-component density profiles (dashed) from the parametric models of Ref. Yang (2024) with bayrons and Ref. Yang et al. (2024) without baryons. The resulting density profiles in SIDM2v are denser than those of the one-component SIDM01 model ($\sigma/m=0.1~\rm cm^{2}/g$), even in the presence of baryons. Since observational constraints depend on the inner mass distribution, our results imply that clusters in SIDM2v are less constrained than those in SIDM01. In both cases, baryons induce a small contraction that does not alter this conclusion.
Appendix E D. Uncertainties in the GGSL cross section
@@
We further quantify the statistical differences in GGSL cross sections among the models. This is done using Welch’s unequal variances $t$-test, which is suitable for comparing the 11 independent realizations per model that may have different variances. The $p$-values and corresponding Gaussian significance in unit of $\sigma$ from this test are presented in Table 3.
@@
Table 3: Statistical significance of GGSL cross section comparisons.
@@
Figure 11: Histograms for the Einstein radii of the contributing secondary critical lines in CDM (blue), Model-2v (red), and Model-x (magenta) models. The corresponding cases that incorporate baryonic effects are shown as dashed lines. A vertical dotted line indicates the veto cut $\theta_{\rm cut}=3^{\prime\prime}$, which is applied to isolate GGSL effects in regions containing observed critical lines. The solid black line shows the observed number density, as taken from Fig. 7 of Ref. Meneghetti et al. (2022).
Figure 12: Convergence test for a $10^{11}~\rm M_{\odot}$ halo with $1\sigma$ concentration (top) and a $10^{13}~\rm M_{\odot}$ halo with median concentration (bottom). For each halo, we consider $N=5\times 10^{4}$ (blue), $10^{5}$ (orange), $5\times 10^{5}$ (red), and $10^{6}$ (magenta) simulation particles. The figure shows the evolution of their inner densities at several enclosed radii.
Appendix F E. Candidates contributing to the GGSL cross sections
Refs. Meneghetti et al. (2022, 2023); Dutra et al. (2025) demonstrate that the subhalos contributing to GGSL predominantly have bound masses in the range $10^{10}-10^{12}~\rm M_{\odot}/h$ and effective Einstein radii $\theta_{E}$ between $0.5^{\prime\prime}$ and $3^{\prime\prime}$. Here, the Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ is the area enclosed by the critical curve.
To confirm that the small-scale lenses in our study are consistent with these findings, we identify subhalos with $\theta_{E}>0.5^{\prime\prime}$ and plot their $\theta_{E}$ distribution in Fig. 11 for the models in Table 2. The number density is obtained by dividing the total number of contributing critical lines by the number of considered systems (11) and by the fiducial calculation area of $100^{\prime\prime}\times 100^{\prime\prime}$. While a few systems have large Einstein radii ($\theta_{E}>3^{\prime\prime}$), the majority of candidates lie in the range $\theta_{E}<3^{\prime\prime}$, in good agreement with observations. Notably, the number densities in our benchmark models, especially those including baryons, show distributions that match the observational data. This agreement further supports the interpretation that SIDM models with mass segregation can produce subhalo populations consistent with GGSL observations.
@@
Appendix G F. Convergence test for halos in the cosmological simulation
This work involves two types of simulations. The controlled simulations are high-resolution, typically using $10^{6}$ or more particles per halo. The cosmological simulations, on the other hand, have a fixed average particle mass of $m_{0}\equiv(m_{H}+m_{L})/2=10^{8}~\rm M_{\odot}$. This resolution implies that a $5\times 10^{12}~\rm M_{\odot}$ halo is resolved with only $5\times 10^{4}$ particles.
@@
Figure 12 presents the results. It shows the mean inner density evolution at $1$, 22, and $3$ kpc for the $10^{11}~\rm M_{\odot}$ halo, and at $4$, 88, and $12$ kpc for the $10^{13}~\rm M_{\odot}$ halo. The tracks for all resolutions converge with the highest-resolution ($N=10^{6}$) result. While simulations with fewer particles show increased fluctuations, their overall evolution trend aligns with the high-resolution result.
Appendix H G. Summary of controlled simulations
We summarize in Table 4 the parameters of the high-resolution controlled simulations in this work.
Simulation $M_{h}({\rm M_{\odot}})$ $\rho_{s}({\rm M_{\odot}/kpc^{3}})$ $r_{s}({\rm kpc})$ $M_{s}({\rm M_{\odot}})$ $r_{H}({\rm kpc})$ SIDM model $N_{\rm part}$
@@
Table 4: Parameters of the controlled N-body simulation halos. The labels “p1”, “p2”, and “m1” in the benchmark names refer to simulations with concentrations of $1\sigma$, $2\sigma$, and $1\sigma$ relative to the median. The labels “B” imply the presence of baryons. Halo masses are determined as the mass enclosed within radii where the density reaches 200 times the critical density ($R_{200}$), based on the scale density ($\rho_{s}$) and scale radius ($r_{s}$) of the NFW profile. The total halo masses in the simulations are slightly higher due to the smooth truncation of the density profiles. Stellar masses are calculated as total masses using the scale density ($\rho_{H}$) and scale radius ($r_{H}$) of the Hernquist profile. Dark matter and stellar simulation particles have equal mass. $N_{\rm part}$ denotes a summation of their numbers.

### v3

@@ A.3 Incorporating the influence of baryons
where $\rho_{\rm s}$, and $r_{\rm s}$ are the NFW scale parameters and $\rho_{\rm H}$, and $r_{\rm H}$ are the Hernquist scale parameters. The dimensionless quantities are defined as $\hat{\rho}_{H}=\rho_{\rm H}/\rho_{\rm s}$, $\hat{r}_{H}=r_{\rm H}/r_{\rm s}$, and $\hat{r}_{\rm eff}=r_{\rm eff}/r_{\rm s}=(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{3}/2)/(1+1.6\hat{\rho}_{H}\hat{r}_{H}^{2}/2)$. To obtain predictions that incorporate baryonic effects, we simply multiply this factor by the core size $r_{c}(\tau)$ in the parametric model for DM-only halos. While schematic, this method accounts for the baryon-induced contraction effect, which scales with the stellar-to-halo mass ratio, and smoothly reduces to the no-baryon limit as $\tau$ approaches zero.
In the main text, we have illustrate the simulation results for a median concentration halo in Fig. 3a, where baryons have growing cores along with the dark matter. The parametric model predictions are tested against the simulation results at different times. Here we provide another example for a more compact halo. In Fig. S2c, we compare the parametric model predictions with the simulation results for a $M^{11}~\rm M_{\odot}$ halo ($1\sigma$ concentration). The profiles, shown in progressively darker colors, correspond to $t=0,3,6,$ and $9$ Gyr, respectively. While small deviations in shape are present, the model effectively captures the growth of inner densities in the baryonic case (solid lines) compared to the DM-only cases (dashed lines). The baryon profiles themselves (dotted lines) remain largely unchanged in this benchmark.
Appendix B B. Calculation of strong lensing effects
[…]
Fig. S3: Density profiles of cluster subhalos with $M>10^{11}\,M_{\odot}/h$ in CDM (a), SIDM2c (b), and SIDM2c+baryons (c) from parametric model predictions (dashed). Solid curves show corresponding fits using the PSIDM model of Refs. Hou et al. (2025). Subpanels display the relative deviation between simulation and fit, defined as $2\,(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The parametric model accurately captures the density structure of simulated subhalos across a broad mass range.
@@
Specifically, the parametric SIDM (PSIDM) density profile introduced in Ref. Hou et al. (2025) has the same parametric form as Eq. (S2). The evolution of its halo parameters is governed by the functions given in that reference. For CDM halos, a single PSIDM profile is sufficient for an accurate fit, as it reduces to the NFW profile in the limit $\tau\to 0$. For SIDM2c halos, we use a superposition of two PSIDM profiles to achieve an accurate representation.
Fig. S3 shows the fitting performance for cluster subhalos in three models: from left to right, CDM (blue), SIDM2c (red), and SIDM2c+baryons (magenta). Only halos with masses above $10^{11}~\rm M_{\odot}$ are shown. The lower panels illustrate the relative difference between the simulated (Sim) and fitted (Fit) profiles, computed as $2(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. Clearly, the two PSIDM model reproduces the simulated halo profiles with high accuracy, even in the presence of baryonic effects.
@@
Appendix C C. Cluster density profiles and constraint
[…]
Fig. S4: (a): Density profiles of cluster host halos in the CDM, SIDM2v, and SIDMx simulations (dashed). The profiles are fitted by two PSIDM density profiles (solid) Hou et al. (2025). (b): Comparison of cluster density profiles at $t=5~\rm Gyr$ for the two-component SIDM2v model (solid) and the one-component SIDM01 model (dashed). The initial CDM halo (black) and baryons (orange) are also shown. The inner densities for SIDM01 are consistently shallower than those for SIDM2v, both with and without baryons, indicating that SIDM2v remains consistent with cluster-scale constraints.
The reconstruction of cluster mass profiles has presented one of the most robust constraints on SIDM in clusters, which varies but is generically around $0.1~\rm cm^{2}/g$ Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). Core formation in clusters is efficient and only moderately influenced by baryon-induced contraction. For a constant, small self-interaction cross section per unit mass, the rate of gravothermal evolution, quantified by the inverse of the core-collapse time, is proportional to $\rho_{\rm s}^{3/2}r_{\rm s}$, implying that more massive halos evolve more rapidly. Moreover, the stellar mass in clusters generally constitutes only about $0.05\%$ of the total halo mass, and thus contributes negligibly to contraction.
@@
Fig. S4a presents the density profiles of host halos from cosmological simulations. As a result of the merging environment and its interplay with inter- and intra-species scattering, we observe intriguing twists in the density profiles of both SIDM2v and SIDMx, which complicates the assessment of cluster-scale constraints.
To obtain a robust assessment, we perform isolated simulations for the cluster halo under SIDM2v, using the scale parameters of the CDM cluster halo as a reference. We consider two contrasting scenarios with and without baryons. The baryonic component, with a mass of $7.47\times 10^{11}~\rm M_{\odot}$, is modeled as a Hernquist profile with $r_{\rm H}=34.1~\rm kpc$. These parameters align with the median stellar-to-halo mass relation and median size-mass relation Moster et al. (2013); Carleton et al. (2019). We use $5\times 10^{6}$ dark matter simulation particles in both cases to better resolve the inner halo regions. Baryons are treated as collisionless stars with particle masses equal to those of the dark matter.
Fig. S4b shows the simulation results (solid) at $t=5~\rm Gyr$, compared with the one-component density profiles (dashed) from the parametric models of Ref. Yang (2024) with bayrons and Ref. Yang et al. (2024) without baryons. The resulting density profiles in SIDM2v are denser than those of the one-component SIDM01 model ($\sigma/m=0.1~\rm cm^{2}/g$), even in the presence of baryons. Since observational constraints depend on the inner mass distribution, our results imply that clusters in SIDM2v are less constrained than those in SIDM01. In both cases, baryons induce a small contraction that does not alter this conclusion.
Appendix D D. Uncertainties in the GGSL cross section
@@
We further quantify the statistical differences in GGSL cross sections among the models. This is done using Welch’s unequal variances $t$-test, which is suitable for comparing the 11 independent realizations per model that may have different variances. The $p$-values and corresponding Gaussian significance in unit of $\sigma$ from this test are presented in Table S1.
@@
Table S1: Statistical significance of GGSL cross section comparisons.
@@
Fig. S5: Histograms for the Einstein radii of the contributing secondary critical lines in CDM (blue), Model-2v (red), and Model-x (magenta) models. The corresponding cases that incorporate baryonic effects are shown as dashed lines. A vertical dotted line indicates the veto cut $\theta_{\rm cut}=3^{\prime\prime}$, which is applied to isolate GGSL effects in regions containing observed critical lines. The solid black line shows the observed number density, as taken from Fig. 7 of Ref. Meneghetti et al. (2022). This profile-level comparison demonstrates that SIDM with mass segregation contributes candidates in the range required to explain the GGSL observations.
[…]
Fig. S6: Convergence test for a $10^{11}~\rm M_{\odot}$ halo with $1\sigma$ concentration (top) and a $10^{13}~\rm M_{\odot}$ halo with median concentration (bottom). For each halo, we consider $N=5\times 10^{4}$ (blue), $10^{5}$ (orange), $5\times 10^{5}$ (red), and $10^{6}$ (magenta) simulation particles. The figure shows the evolution of their inner densities at several enclosed radii.
Appendix E E. Candidates contributing to the GGSL cross sections
This section provides a profile-level check that complements the total cross-section comparison presented in the main text and the previous appendix. Refs. Meneghetti et al. (2022, 2023); Dutra et al. (2025) demonstrate that the subhalos contributing to GGSL predominantly have bound masses in the range $10^{10}-10^{12}~\rm M_{\odot}/h$ and effective Einstein radii $\theta_{E}$ between $0.5^{\prime\prime}$ and $3^{\prime\prime}$. Here, the Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ is the area enclosed by the critical curve.
To confirm that the small-scale lenses in our study are consistent with these findings, we identify subhalos with $\theta_{E}>0.5^{\prime\prime}$ and plot their $\theta_{E}$ distribution in Fig. S5 for the models in Table 2. The number density is obtained by dividing the total number of contributing critical lines by the number of considered systems (11) and by the fiducial calculation area of $100^{\prime\prime}\times 100^{\prime\prime}$. While a few systems have large Einstein radii ($\theta_{E}>3^{\prime\prime}$), the majority of candidates lie in the range $\theta_{E}<3^{\prime\prime}$, in good agreement with observations. Notably, the number densities in our benchmark models, especially those including baryons, show distributions that match the observational data. This agreement further supports the interpretation that SIDM models with mass segregation can produce subhalo populations consistent with GGSL observations.
@@
Appendix F F. Convergence test for halos in the cosmological simulation
This work involves two types of simulations. The controlled simulations are high-resolution, typically using $10^{6}$ or more particles per halo. The cosmological simulations, on the other hand, have a fixed average particle mass of $m_{0}\equiv(m_{\rm H}+m_{\rm L})/2=10^{8}~\rm M_{\odot}$. This resolution implies that a $5\times 10^{12}~\rm M_{\odot}$ halo is resolved with only $5\times 10^{4}$ particles.
@@
Fig. S6 presents the results. It shows the mean inner density evolution at $1$, 22, and $3$ kpc for the $10^{11}~\rm M_{\odot}$ halo, and at $4$, 88, and $12$ kpc for the $10^{13}~\rm M_{\odot}$ halo. The tracks for all resolutions converge with the highest-resolution ($N=10^{6}$) result. While simulations with fewer particles show increased fluctuations, their overall evolution trend aligns with the high-resolution result.
Table S2: Parameters of the controlled N-body simulation halos.
Simulation $M_{\rm h}({\rm M_{\odot}})$ $\rho_{\rm s}({\rm M_{\odot}/kpc^{3}})$ $r_{\rm s}({\rm kpc})$ $M_{s}({\rm M_{\odot}})$ $r_{\rm H}({\rm kpc})$ SIDM model $N_{\rm part}$
@@
The labels “p1”, “p2”, and “m1” in the benchmark names refer to simulations with concentrations of $1\sigma$, $2\sigma$, and $1\sigma$ relative to the median. The labels “B” imply the presence of baryons. Halo masses are determined as the mass enclosed within radii where the density reaches 200 times the critical density ($R_{200}$), based on the scale density ($\rho_{\rm s}$) and scale radius ($r_{\rm s}$) of the NFW profile. The total halo masses in the simulations are slightly higher due to the smooth truncation of the density profiles. Stellar masses are calculated as total masses using the scale density ($\rho_{\rm H}$) and scale radius ($r_{\rm H}$) of the Hernquist profile. Dark matter and stellar simulation particles have equal mass. $N_{\rm part}$ denotes a summation of their numbers.
Appendix G G. Summary of controlled simulations
We summarize in Table S2 the parameters of the high-resolution controlled simulations in this work.